OpenStudy (sstradic):

The leg of a right triangle is 5 units and the hypotenuse is 8 units. What is the length, in units, of the other leg of the triangle?

OpenStudy (steve816):

Use the Pythagorean Theorem!\[\large a^2+b^2=c^2\]where c is the hypotenuse

OpenStudy (steve816):

So basically, solve this equation for x:\[\large x^2+5^2=8^2\]

OpenStudy (steve816):

Got it?

OpenStudy (sstradic):

not really :/ i have trouble understanding this kinda stuff

OpenStudy (steve816):

Which part do you not get?

OpenStudy (sstradic):

all of it XD

OpenStudy (steve816):

well, basically what the Pythagorean theorem is saying is that when you square the two legs of the triangle and add it up, it should equal to the square of the hypotenuse.

OpenStudy (sstradic):

SO RADICAL MARK THINGY THEN THE NUMBER 39 :3

OpenStudy (steve816):

lol wut??

OpenStudy (sstradic):

a radical then the number 39????????????????????????????/

OpenStudy (steve816):

ohh I got you. Yeah, the answer would be\[\large \sqrt{39}\]

OpenStudy (sstradic):

yay im not retarted lel #freelilkodak

OpenStudy (steve816):

xD gj